OpenStudy (anonymous):

Solve: The formula for degrees Celsius is: C = 5/9 (F − 32), where F stands for degrees Fahrenheit. then Determine how many degrees Fahrenheit 20 degrees Celsius is.

OpenStudy (anonymous):

What are you trying to solve for? F? C = (5/9)(F - 32) (F - 32) = C/(5/9) F = (5/9)C + 32 F = 1.8C + 32

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then Determine how many degrees Fahrenheit 20 degrees Celsius is

OpenStudy (anonymous):

Just plug it into the above formula for F.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

C=5/9(20-32) C=5/9(-12) C=-60/9 C=-20/3